OpenStudy (anonymous):

Definite integral help please. Can someone please help me. I have x^2(x+2) dx from -2 to 2. To start off with I put u = (x+2) du/dx=x and du/x = dx

OpenStudy (mimi_x3):

\[\int\limits x^{2}(x+2)dx \] Just expand it; i don't think u-sub would work.

OpenStudy (anonymous):

I wondered that originally because when I tried to u sub it it didn't work lol. So x^3 + 2x^2

OpenStudy (anonymous):

\[x^3+2x^2 = x^3+x^4/2\]

OpenStudy (mimi_x3):

I don't think so..

OpenStudy (anonymous):

oh can you please show me the way

OpenStudy (mimi_x3):

\[=>\int\limits x^{3}+2x^{2}dx => \frac{x^{3+1}}{3+1} +\frac{2x^{2+1}}{2+1} +C\]

OpenStudy (mimi_x3):

\[=> \frac{x^{4}}{4} +\frac{2x^{3}}{3} +C\]

OpenStudy (anonymous):

Thank you

OpenStudy (mimi_x3):

np